OpenStudy (anonymous):

integral(1/(xsqrt(1+ln(x)^2)dx

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ x \sqrt{1+\ln^2} }dx\]

OpenStudy (anonymous):

u=ln(x) du=1/xdx dx=xdu

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ \sqrt{1+u^2}}du\]

OpenStudy (anonymous):

we can't use the identity so do I have to do a trig sub for this?

OpenStudy (anonymous):

u=sin(theta) dx=cos(theta)

OpenStudy (anonymous):

but when I try this: \[\int\limits\frac{ 1 }{ \sqrt{1+\sin^2}(\theta) }\cos(\theta)d(\theta)\]

OpenStudy (anonymous):

and simplify i get down to \[\int\limits \frac{ 1 }{ \cos(\theta) }\cos(\theta)d(\theta)\] and \[\int\limits1d(\theta)\]

OpenStudy (anonymous):

i get stuck

OpenStudy (anonymous):

what is arcsinh?

OpenStudy (bradely):

inverse hyperbolic function and arcsinhx = log (x+sqrt(x^2+1)) Source: http://www.mathskey.com/question2answer/